OpenStudy (anonymous):

The hypotenuse of a right triangle is 6 m long. One leg is 1 m longer than the other. Find the length of the shorter leg. Round to the nearest hundredth.

OpenStudy (anonymous):

ARMYRANGERS

OpenStudy (anonymous):

@ARMYRANGERS

OpenStudy (anonymous):

@anonymous_user

OpenStudy (anonymous):

@stabar

OpenStudy (armyrangers):

Ok you will have to use the \(\large a^2+b^2=c^2\)

OpenStudy (anonymous):

im not sure what goes were

OpenStudy (armyrangers):

Ok so you know c

OpenStudy (anonymous):

yea

OpenStudy (armyrangers):

Ok so you are going to do Let a = shorter leg b = a + 1 = longer leg a² + b² = 6² a² + (a + 1)² = 6² a² + a² + 2a + 1 = 36 2a² + 2a – 35 = 0 a = {–(2) ± √[(2)² – 4(2)(–35)]/2(2) a = (–2 ± √284)/4 a = (–2 ± 2√71)/2 a = (–1 ± √71)/2 a = (–1 – √71)/2 .................. discard, negative a = (–1 + √71)/2 = 3.7130748865881793153170699531014 a = 3.71 m..................... shorter side b = 4.71 m..................... longer side